Brooks Sporting Inc. is prepared to report the following 2019 income statement (shown in thousands of dollars).

Sales $19,000
Operating costs including depreciation 13,680
EBIT $5,320
Interest 264
EBT $5,056
Taxes (25%) 1,264
Net income $3,792

Prior to reporting this income statement, the company wants to determine its annual dividend. The company has 410,000 shares of common stock outstanding, and its stock trades at $52 per share.

  1. The company had a 55% dividend payout ratio in 2018. If Brooks wants to maintain this payout ratio in 2019, what will be its per-share dividend in 2019? Do not round intermediate calculations. Round your answer to the nearest cent.
    $  

  2. If the company maintains this 55% payout ratio, what will be the current dividend yield on the company's stock? Do not round intermediate calculations. Round your answer to two decimal places.
      %

  3. The company reported net income of $3.5 million in 2018. Assume that the number of shares outstanding has remained constant. What was the company's per-share dividend in 2018? Do not round intermediate calculations. Round your answer to the nearest cent.
    $  

  4. As an alternative to maintaining the same dividend payout ratio, Brooks is considering maintaining the same per-share dividend in 2019 that it paid in 2018. If it chooses this policy, what will be the company's dividend payout ratio in 2019? Do not round intermediate calculations. Round your answer to two decimal places.
      %

  5. Assume that the company is interested in dramatically expanding its operations and that this expansion will require significant amounts of capital. The company would like to avoid transactions costs involved in issuing new equity. Given this scenario, would it make more sense for the company to maintain a constant dividend payout ratio or to maintain the same per-share dividend?
    1. Since the company would like to avoid transactions costs involved in issuing new equity, it would be best for the firm to maintain the same per-share dividend.
    2. Since the company would like to avoid transactions costs involved in issuing new equity, it would be best for the firm to maintain a constant dividend payout ratio.
Answer #1

a.Dividend in 2019 = Net Income*Payout ratio

= 3,792*55%

= $2,085.6

Per share dividend = Total Dividend/Number of shares

= 2,085.6/410

= $5.0868

i.e. $5.09

b.Dividend yield = Dividend per share/Market price per share

= 5.09/52

= 9.79%

c.Per share dividend in 2018 = 3,500,000*55%/410,000

= $4.70

d.Payout ratio = 4.70*410,000/3,792,000

= 50.82%

e.I.same per share dividend as it is lower